\begin{thebibliography}{}
%
% and use \bibitem to create references. Consult the Instructions
% for authors for reference list style.
%

%
\bibitem {pdg2008} C.~Amsler et al. (Particle Data Group), Phys.\ Lett.\ \textbf{B667}, 1
(2008).
%
%\cite{Arndt:2006bf}
\bibitem{Arndt:2006bf}
 R.~A.~Arndt, W.~J.~Briscoe, I.~I.~Strakovsky and R.~L.~Workman,
 ``Extended Partial-Wave Analysis of $\pi N$ Scattering Data,''
 Phys.\ Rev.\  C {\bf 74}, 045205 (2006).
% [arXiv:nucl-th/0605082].
 %%CITATION = PHRVA,C74,045205;%%

%\cite{Arndt:2008zz}
\bibitem{Arndt:2008zz}
 R.~Arndt, W.~Briscoe, I.~Strakovsky and R.~Workman,
 ``Partial-wave analysis and baryon spectroscopy,''
 Eur.\ Phys.\ J.\  A {\bf 35}, 311 (2008).
 %%CITATION = EPHJA,A35,311;%%

%\cite{Arndt:2009nv}
\bibitem{Arndt:2009nv}
 R.~A.~Arndt, W.~J.~Briscoe, M.~W.~Paris, I.~I.~Strakovsky and R.~L.~Workman,
 ``Baryon Resonance Analysis from SAID,''
 Chin.\ Phys.\  C {\bf 33}, 1063 (2009).
% [arXiv:0906.3709 [nucl-th]].
 %%CITATION = CHPHD,C33,1063;%%
%
%\cite{Burkert:2004sk}
\bibitem{Burkert:2004sk}
 V.~D.~Burkert and T.~S.~H.~Lee,
 ``Electromagnetic meson production in the nucleon resonance region,''
 Int.\ J.\ Mod.\ Phys.\  E {\bf 13}, 1035 (2004).
% [arXiv:nucl-ex/0407020].
%\cite{Matsuyama:2006rp}
\bibitem{Matsuyama:2006rp}
 A.~Matsuyama, T.~Sato and T.~S.~Lee,
 ``Dynamical coupled-channel model of meson production reactions in the
 nucleon resonance region,''
 Phys.\ Rept.\  {\bf 439}, 193 (2007).
% [arXiv:nucl-th/0608051].
 %%CITATION = PRPLC,439,193;%%

%\cite{JuliaDiaz:2007kz}
\bibitem{JuliaDiaz:2007kz}
 B.~Julia-Diaz, T.~S.~Lee, A.~Matsuyama and T.~Sato,
 ``Dynamical Coupled-Channel Model of $\pi N$ Scattering in the W $\leq$ 2 GeV
 Nucleon Resonance Region,''
 Phys.\ Rev.\  C {\bf 76}, 065201 (2007).
% [arXiv:0704.1615 [nucl-th]].
 %%CITATION = PHRVA,C76,065201;%%

%\cite{Anisovich:2009zy}
\bibitem{Anisovich:2009zy}
 A.~V.~Anisovich, E.~Klempt, V.~A.~Nikonov, M.~A.~Matveev, A.~V.~Sarantsev and U.~Thoma,
 ``Photoproduction of pions and properties of baryon resonances from a
 Bonn-Gatchina partial wave analysis,''
 Eur.\ Phys.\ J.\  A {\bf 44}, 203 (2010).
% [arXiv:0911.5277 [hep-ph]].
 %%CITATION = EPHJA,A44,203;%%

%\cite{Horn:2007pp}
\bibitem{Horn:2007pp}
 I.~Horn {\it et al.}  [CB-ELSA Collaboration],
 ``Evidence for a parity doublet $\Delta(1920)P_{33}$ and $\Delta(1940)D_{33}$
 from $\gamma p\to p\pi^0\eta$,''
 Phys.\ Rev.\ Lett.\  {\bf 101}, 202002 (2008).
% [arXiv:0711.1138 [nucl-ex]].
 %%CITATION = PRLTA,101,202002;%%

%\cite{Klempt:2006sa}
\bibitem{Klempt:2006sa}
 E.~Klempt, A.~V.~Anisovich, V.~A.~Nikonov, A.~V.~Sarantsev and U.~Thoma,
 ``Phase Motion Of Baryon Resonances,''
 Eur.\ Phys.\ J.\  A {\bf 29}, 307 (2006).
 %%CITATION = EPHJA,A29,307;%%

%\cite{Arends:2008zz}
\bibitem{Arends:2008zz}
 H.~J.~Arends,
 ``Recent activities and perspectives at MAMI,''
 AIP Conf.\ Proc.\  {\bf 1056}, 428 (2008).
 %%CITATION = APCPC,1056,428;%%

%
%\cite{MartinezTorres:2007sr}
\bibitem{MartinezTorres:2007sr}
 A.~Martinez Torres, K.~P.~Khemchandani and E.~Oset,
 ``Three body resonances in two meson-one baryon systems,''
 Phys.\ Rev.\  C {\bf 77}, 042203 (2008).
% [arXiv:0706.2330 [nucl-th]].
 %%CITATION = PHRVA,C77,042203;%%

%\cite{Khemchandani:2008rk}
\bibitem{Khemchandani:2008rk}
 K.~P.~Khemchandani, A.~Martinez Torres and E.~Oset,
 ``The $N^*(1710)$ as a resonance in the $\pi\pi N$ system,''
 Eur.\ Phys.\ J.\  A {\bf 37}, 233 (2008).
% [arXiv:0804.4670 [nucl-th]].
 %%CITATION = EPHJA,A37,233;%%


%\cite{MartinezTorres:2009cw}
\bibitem{meissner}
 A.~Martinez Torres, K.~P.~Khemchandani, U.~G.~Meissner and E.~Oset,
 ``Searching for signatures around 1920-MeV of a $N^*$ state of three hadron
 nature,''
 Eur.\ Phys.\ J.\  A {\bf 41}, 361 (2009).
% [arXiv:0902.3633 [nucl-th]].
 %%CITATION = EPHJA,A41,361;%%
%
%\cite{Jido:2008kp}
\bibitem{Jido:2008kp}
 D.~Jido and Y.~Kanada-En'yo,
 ``K anti-K N molecule state with $I=1/2$ and $J^P=1/2^+$ studied with three-body
 calculation,''
 Phys.\ Rev.\  C {\bf 78}, 035203 (2008).
% [arXiv:0806.3601 [nucl-th]].
 %%CITATION = PHRVA,C78,035203;%%


%\cite{MartinezTorres:2008kh}
\bibitem{MartinezTorres:2008kh}
 A.~Martinez Torres, K.~P.~Khemchandani and E.~Oset,
 ``Solution to Faddeev equations with two-body experimental amplitudes as
 input and application to $J^P=1/2^+$, S=0 baryon resonances,''
 Phys.\ Rev.\  C {\bf 79}, 065207 (2009).
% [arXiv:0812.2235 [nucl-th]].
 %%CITATION = PHRVA,C79,065207;%%

%\cite{Chand:1962ec}
\bibitem{Chand:1962ec}
 R.~Chand and R.~H.~Dalitz,
 ``Charge-independence in $K^-d$ capture reactions,''
 Annals Phys.\  {\bf 20}, 1 (1962)
 %%CITATION = APNYA,20,1;%%

  %\cite{Barrett:1999cw}
\bibitem{Barrett:1999cw}
 R.~C.~Barrett and A.~Deloff,
 ``Strong interaction effects in kaonic deuterium,''
 Phys.\ Rev.\  C {\bf 60}, 025201 (1999).
 %%CITATION = PHRVA,C60,025201;%%
  %\cite{Deloff:1999gc}
\bibitem{Deloff:1999gc}
 A.~Deloff,
 ``$\eta d$ and $K^- d$ zero energy scattering: A Faddeev approach,''
 Phys.\ Rev.\  C {\bf 61}, 024004 (2000).
 %%CITATION = PHRVA,C61,024004;%%


 %\cite{Kamalov:2000iy}
\bibitem{Kamalov:2000iy}
 S.~S.~Kamalov, E.~Oset and A.~Ramos,
 ``Chiral unitary approach to the $K^- d$ scattering length,''
 Nucl.\ Phys.\  A {\bf 690}, 494 (2001).
%  [arXiv:nucl-th/0010054].
 %%CITATION = NUPHA,A690,494;%%

%\cite{Roca:2010tf}
\bibitem{Roca:2010tf}
  L.~Roca and E.~Oset,
  ``A description of the $f_2(1270), \rho_3(1690), f_4(2050), \rho_5(2350)$ and
  $f_6(2510)$ resonances as multi-$\rho(770)$ states,''
  Phys.\ Rev.\  D {\bf 82}, 054013 (2010).
%  [arXiv:1005.0283 [hep-ph]].
  %%CITATION = PHRVA,D82,054013;%%

%\cite{YamagataSekihara:2010qk}
\bibitem{YamagataSekihara:2010qk}
  J.~Yamagata-Sekihara, L.~Roca and E.~Oset,
  ``On the nature of the $K^*_2(1430), K^*_3(1780), K^*_4(2045), K^*_5(2380)$ and $K^*_6$ as
  $K^*$--multi-$\rho$ states,''
arXiv:1010.0525 [hep-ph], Phys. Rev. D, in print.
  %%CITATION = ARXIV:1010.0525;%%
%

%
\bibitem{vrana00}T.~P.~Vrana, S.~A.~Dytman, and T.~S.~H.~Lee, ``Baryon Resonance Extraction from piN Data using a Unitary Multichannel
  Model,''
 Phys.\ Rept.\ \textbf{328}, 181
(2000).
%
\bibitem{manley92}D.~M.~Manley, and E.~M.~Saleski, ``Multichannel Resonance Parametrization Of $\pi N$ Scattering Amplitudes,''
 Phys.\ Rev.\ D
\textbf{45}, 4002 (1992).
%
\bibitem{cutkosky80}R.~E.~Cutkosky et al., Presented at 4th Int. Conf. on Baryon Resonances, Toronto, Canada, Jul 14-16, 1980.
Published in Baryon 1980:19 (QCD161:C45:1980).


\end{thebibliography}

\documentclass{svjour3}                     % onecolumn (standard format)
\smartqed  % flush right qed marks, e.g. at end of proof
\usepackage{graphicx}
\begin{document}

\title{$\Delta \rho \pi$ interaction leading to $N^*$ and $\Delta^*$ resonances
\footnote{Presented at the 21st European Conference on Few-Body
Problems in Physics, Salamanca, Spain, 30 August - 3 September
2010.} }

%\titlerunning{Short form of title}        % if too long for running head

\author{Ju-Jun Xie \and A. Mart\'inez Torres  \and E. Oset \and P. Gonz\'alez.
}

\institute{Ju-Jun Xie \at Instituto de F\'\i sica Corpuscular
(IFIC), Centro Mixto CSIC-Universidad de Valencia, Institutos de
Investigaci\'on de Paterna, Aptd. 22085, E-46071 Valencia, Spain \\
Department of Physics, Zhengzhou University, Zhengzhou,
Henan 450001, China \\ \email{xiejujun@ific.uv.es}           %  \\
%             \emph{Present address:} of F. Author
%  if neede
\and A. Mart\'inez Torres \at Yukawa Institute for Theoretical
Physics, Kyoto University, Kyoto 606-8502, Japan \and E. Oset and P.
Gonz\'alez \at Departamento de F\'\i sica Te\'orica and Instituto de
F\'\i sica Corpuscular (IFIC), Centro Mixto CSIC-Universidad de
Valencia, Institutos de Investigaci\'on de Paterna, Aptd. 22085,
E-46071 Valencia, Spain
%\and P. Gonz\'alez \at Departamento de F\'\i sica Te\'orica and
%Instituto de F\'\i sica Corpuscular (IFIC), Centro Mixto
%CSIC-Universidad de Valencia, Institutos de Investigaci\'on de
%Paterna, Aptd. 22085, E-46071 Valencia, Spain
 }

\date{Received: date / Accepted: date}
% The correct dates will be entered by the editor

\maketitle

\begin{abstract}
We have performed a calculation for the three body $\Delta \rho \pi$
system by using the fixed center approximation to Faddeev equations,
taking the interaction between $\Delta$ and $\rho$, $\Delta$ and
$\pi$, and $\rho$ and $\pi$ from the chiral unitary approach. We
find several peaks in the modulus squared of the three-body
scattering amplitude, indicating the existence of resonances, which
can be associated to known $I=1/2, 3/2$ and $J^P=1/2^+, 3/2^+$ and
$5/2^+$ baryon states.

\keywords{Fixed center approximation \and Three body system \and
Chiral unitary model}
% \PACS{PACS code1 \and PACS code2 \and more}
% \subclass{MSC code1 \and MSC code2 \and more}
\end{abstract}

\section{Introduction}
\label{intro}

Our knowledge on the baryon resonances mainly comes from $\pi N$
experiments and is still under
debate~\cite{pdg2008,Arndt:2006bf,Arndt:2008zz,Arndt:2009nv}. The
information extracted from photon nucleon reactions have helped in
making progress in this field, reconforming many known resonances
and claiming evidence for new
ones~\cite{Burkert:2004sk,Matsuyama:2006rp,JuliaDiaz:2007kz,Anisovich:2009zy,Horn:2007pp,Klempt:2006sa,Arends:2008zz}.
The fact that some known resonances are explained in terms of three
body systems of two mesons and one
baryon~\cite{MartinezTorres:2007sr,Khemchandani:2008rk} should
certainly stimulate work looking for resonances in three body final
states of reactions. In this sense a suggestion is made
in~\cite{meissner} to look for a predicted state of
$N\bar{K}K$~\cite{Jido:2008kp,MartinezTorres:2008kh} in the $\gamma
p \to K^+ K^- p$ reaction close to threshold.

The main aim of the present work is to investigate the three-body
$\Delta \rho \pi$ system considering the interaction of the three
components among themselves keeping in mind the expected strong
correlations of the $\Delta \rho$ system which generate the $N^*$
and/or $\Delta^*$ bound states. For this purpose, we have solved the
Faddeev equations by using the fixed center approximation(FCA) in
terms of two body $\Delta \pi$ and $\rho \pi$ scattering amplitudes.

The FCA to the Faddeev equations is a tool which has proved to be
efficient and accurate to study the interaction of particles with
bound states of a pair of particles at very low energies, or below
threshold~\cite{Chand:1962ec,Barrett:1999cw,Deloff:1999gc,Kamalov:2000iy}.
Recently, this approach was used in Ref.~\cite{Roca:2010tf} to
describe the $f_2(1270)$, $\rho_3(1690)$, $f_4(2050)$,
$\rho_5(2350)$ and $f_6(2510)$ resonances as multi-$\rho(770)$
states, and in Ref.~\cite{YamagataSekihara:2010qk} to study the
$K^*_2(1430)$, $K_3^*(1780)$, $K^*_4(2045)$, $K^*_5(2380)$, and a
not yet discovered $K^*_6$ resonances as $K^*-$multi$-\rho$ states.
The success of these works encourages us to extend the method to
study the present $\Delta \rho \pi$ system.

\section{Formalism and results}

For the three body $\Delta \rho \pi$ system, we consider $\Delta
\rho$ as a bound state of $N^*$($I_{\Delta \rho}=1/2$) resonance or
$\Delta^*$($I_{\Delta \rho}=3/2$) resonance, which allows us to use
the FCA to solve the Faddeev equations. The external $\pi$ meson
interacts successively with the $\Delta$ baryon and the $\rho$ meson
which form the $\Delta \rho$ cluster. In terms of two partition
functions $T_1$ and $T_2$, the FCA equations are
\begin{eqnarray}
T_1 &=& t_1+t_1G_0T_2 , \label{T1} \\
T_2 &=& t_2+t_2G_0T_1, \label{T2} \\
T &=& T_1+T_2,        \label{T}
\end{eqnarray}
where $T$ is the total three-body scattering amplitude and
$T_i$($i=1,2$) accounts for the diagrams starting with the
interaction of the external particle with the particle $i$ of the
compound system and $t_i$ represents the two body $\Delta \pi$ and
$\rho \pi$ unitarized scattering amplitudes.

Next, we will show the results obtained from the scattering
amplitude of the $\Delta \rho \pi$ system. We evaluate the
scattering amplitude $T$ matrix of Eq.~(\ref{T}) and associate the
peaks of $|T|^2$ to resonances. In table~\ref{tab:results} we show a
summary of the findings obtained from our model and the tentative
association to known states~\cite{pdg2008}.

\begin{table}[htbp]
\caption{The properties of the generated resonances with our model
and their possible PDG counterparts. \label{tab:results}}
\begin{center}
\begin{tabular}{c|c|cccc}
\hline  $I_{\Delta \rho}, I_{total}$ & Mass of our  & \multicolumn{4}{c}{PDG data} \\
\vspace*{-0.3cm}
       &  &  & & &   \\
      & model(MeV)  & name &$J^P$  & mass(MeV)  & status \\
\hline $\frac{1}{2},\frac{1}{2}$ & $\sim 1850$
& $N^*(1900)$ & $3/2^+$ & $  1900$ & ** \\
\hline $\frac{1}{2},\frac{3}{2}$ & $\sim 1800$ &
 $\Delta^*(1750)$ & $1/2^+$ & $ 1750$ & *   \\
 &  &$\Delta^*(2000)^?$ & $5/2^+$ & $1724 \pm 61$ & Ref.~\cite{vrana00}  \\
 &  &$\Delta^*(2000)^?$ & $5/2^+$ & $1752 \pm 32$ & Ref.~\cite{manley92}  \\
 &$\sim 1900$  &$\Delta^*(1905)$ & $5/2^+$ & $1865 - 1915$ & ****  \\
 &  &$\Delta^*(1920)$ & $3/2^+$ & $1900 - 1970$ & ***  \\
 &$\sim 2200$  &$\Delta^*(2000)^?$ & $5/2^+$ & $2200 \pm 125$ & Ref.~\cite{cutkosky80}  \\
\hline $\frac{3}{2},\frac{1}{2}$ & $\sim 2000$
& $N^*(2000)$ & $5/2^+$ & $  2000$ &** \\
\hline $\frac{3}{2},\frac{3}{2}$ & $\sim 2000$
& ? &?  &? &? \\
\hline

\end{tabular}
\end{center}
\end{table}

\section{Discussions and Conclusions}

We have performed a Faddeev calculation for the three body $\Delta
\rho \pi$ system by using the fixed center approximation, taking the
interaction between $\Delta$ and $\rho$, $\Delta$ and $\pi$, and
$\rho$ and $\pi$ from the chiral unitary approach. The $\Delta \rho$
interaction within the framework of the hidden-gauge formalism in
$I=1/2$ sector describes the $N^*(1675)J^P=5/2^-$ as a $\Delta \rho$
bound state, then we write the three-body interaction in terms of
two-body ($\Delta \pi$ and $\rho \pi$) $s-$wave scattering
amplitudes based on the chiral Lagrangians. The three body states
found are degenerated in $J^P=1/2^+,3/2^+,5/2^+$. We found
candidates in the PDG book which can be associated to the states
obtained, but one of them, with isospin $3/2$ and mass around $2000$
MeV, is missing. It is very interesting to observe that, even if the
$\Delta \rho \pi$ system allows for $I=5/2$, the dynamics of the
system precludes the formation of these exotic states.

\bigskip
\noindent

\begin{acknowledgements}
This work is partly supported by DGICYT Contract No. FIS2006-03438,
the Generalitat Valenciana in the project PROMETEO and the EU
Integrated Infrastructure Initiative Hadron Physics Project under
contract RII3-CT-2004-506078. Ju-Jun Xie acknowlwdges Ministerio de
Educaci\'on Grant SAB2009-0116. The work of A.~M.~T.~is supported by
the Grant-in-Aid for the Global COE Program ``The Next Generation of
Physics, Spun from Universality and Emergence" from the Ministry of
Education, Culture, Sports, Science and Technology (MEXT) of Japan.
\end{acknowledgements}

%\begin{acknowledgements}
%If you'd like to thank anyone, place your comments here
%and remove the percent signs.
%\end{acknowledgements}

% BibTeX users please use one of
%\bibliographystyle{spbasic}      % basic style, author-year citations
%\bibliographystyle{spmpsci}      % mathematics and physical sciences
%\bibliographystyle{spphys}       % APS-like style for physics
%\bibliography{}   % name your BibTeX data base

% Non-BibTeX users please use

\end{document}